OpenStudy (anonymous):

hi pls find x. i can't seem to factor it. 3x^2 +7x = x -2

terenzreignz (terenzreignz):

This can be simplified to: 3x^2 + 6x + 2 = 0 If all else fails, we can always use the quadratic formula, or: We divide the entire thing by 3: x^2 + 2x + (2/3) = 0 And then, we complete the square: x^2 + 2x + (2/3) + 1 - 1= 0 ---> (we added 1 and then subtracted 1, effectively nullifying) (x^2 + 2x + 1) + (2/3) - 1 = 0 (x + 1)^2 - (1/3) = 0 x + 1 = sqrt(1/3) x = -1 - sqrt(1/3) OR x = -1 + sqrt(1/3)

OpenStudy (anonymous):

\[3x^2+7x=x-2\] \[3x^2+7x-x+2=0\] \[3x^2+6x+2=0\] We can't factorise this so quadratic formula ahoy! Where: a=3 b=6 c=2 \[x=(-b \pm \sqrt{b^2-4ac})/2a\] \[x=(-6\pm \sqrt{6^2-4(3)(2)})/6\] \[x=(-6\pm \sqrt{12})/6\] \[x=(-6\pm2\sqrt{3})/6\] \[x=(-3\pm \sqrt{3})/3\] \[x=-1\pm \sqrt{3}/3\] Hope that helps :)